OpenStudy (anonymous):

Find an equation of a line through (−2, −3) with M=1/2

OpenStudy (solomonzelman):

\(\large\color{black}{ (y-y_1)=m(x-x_1) }\)

OpenStudy (solomonzelman):

for example, if I was asked: `Find an equation of a line through (-3, 2) with M=1/3` the this is what I would do: \(\large\color{black}{ y-y_1=m(x-x_1) }\) \(\large\color{black}{ y-2=\frac{\LARGE 1}{\LARGE3}(x-(-3)~) }\) \(\large\color{black}{ y-2=\frac{\LARGE 1}{\LARGE3}(x+3) }\) \(\large\color{black}{ y-2=\frac{\LARGE 1}{\LARGE3}x+1 }\) \(\large\color{black}{ y=\frac{\LARGE 1}{\LARGE3}x+1 +2 }\) \(\large\color{blue}{ y=\frac{\LARGE 1}{\LARGE3}x+3 }\)

OpenStudy (anonymous):

what did you do on your 3rd step @SolomonZelman

OpenStudy (solomonzelman):

I know that: \(\large\color{black}{ a-(-b)=a+b }\) correct?

OpenStudy (anonymous):

sorry i meant 4th i understood the 3rd one

OpenStudy (solomonzelman):

yes, so I just expanded the parenthesis on the right side: \(\large\color{black}{ \frac{\LARGE 1}{\LARGE 3}(x+3) \\ \\ \Longrightarrow (\frac{\LARGE 1}{\LARGE 3})x+(\frac{\LARGE 1}{\LARGE 3})3 \\ \\ \frac{\LARGE 1}{\LARGE 3}x+1 }\)

OpenStudy (solomonzelman):

makes sense?

OpenStudy (solomonzelman):

if you have any questions regarding your problem, I can surely help you./

OpenStudy (anonymous):

yes thank you

OpenStudy (solomonzelman):

yes, you have questions, or yes, you are good.... ? ;)